OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (3, 6) and (−2, 1)?

OpenStudy (anonymous):

Are there options?

OpenStudy (anonymous):

Hmm. I'd say 1.

OpenStudy (anonymous):

y + 6 = 1(x + 3) y + 1 = −1(x + 2) y − 6 = −1(x + 3) y − 6 = 1(x − 3)

OpenStudy (anonymous):

I'd ask @Baseballguy101 He's good at this stuff.

OpenStudy (baseballguy101):

do you know the formula first of all?

OpenStudy (anonymous):

nope

OpenStudy (baseballguy101):

ok it's y-y1 = m(x-x1). what do you think we shoud do from here?

OpenStudy (anonymous):

6+2=m(1-3) idk

OpenStudy (baseballguy101):

before we go further do you have a graph or is that it?

OpenStudy (anonymous):

that's it

OpenStudy (baseballguy101):

ok so with your equation being y-y1 = m(x-x1), m being slope, we will need to find this. Do you know how?

OpenStudy (zzr0ck3r):

Given two points \((x_1,y_1),(x_2,y_2)\) first find the slope given by \(m=\dfrac{y_1-y_0}{x_1-x_0}\) and then choose any point (say the first one) and we have \[y=m(x-x_1)-y_1\]

OpenStudy (anonymous):

i have nno idea about any of this to be honest

OpenStudy (mathmate):

@nigel305 These sites might help you if you need to review the concepts. They are easy to read and help you recall what you have learned before. http://www.algebra.com/algebra/homework/Graphs/Slope-of-a-Line-between-two-points.lesson https://www.mathsisfun.com/algebra/line-equation-point-slope.html